# of the a an and or to from in as about but than then so through

### v1

@@ 5.2. File-based Analysis
After downloading the repositories and retaining only the files written in statically typed languages with floating-point keywords, we are left with 13 297 30013\,297\,300 files. To estimate the false positive rate, we draw a random sample from this dataset and manually inspect each file to determine whether it contains actual floating-point code. Using Cochran’s formula (Cochran, 1977) for sample size estimation, we select 385 files, ensuring statistical representativeness at a 95 % confidence level with a 5 % margin of error. Of these 385 files, 310 contain valid floating-point code, while 75 are false positives. Among the false positives, 52 arise from keywords appearing in comments, 7 from keywords in string literals, 5 from keywords used in variable names, 3 from misclassified file extensions, and 8 from other miscellaneous causes. In most cases, the keyword double appears in comments in C and C++ files. This indicates the importance of stripping comments and ignoring string literals before performing keyword-based detection at the function level in later steps of the methodology.
Given the results of the manual inspection and using Wilson score intervals (and, 1927), we estimate that the proportion of true positives in our dataset lies between 76 % and 85 %, with 95 % confidence. Assuming that the presence of false positives in files is independent across files, the probability that project $i$ contains at least one file with actual floating-point code is
@@
Deduplication
A substantial fraction of the dataset is redundant, with 62 % of files identified as token-level duplicates and 61 % as exact duplicates. TypeScript exhibits the highest duplication rate, with declaration files duplicated on average 17 times and source files twice. C follows, with 3.4 copies per source file and 2.8 per header file. The most duplicated file originates from Angular, a TypeScript web application framework, and appears 27 65227\,652 times (0.7 % of TypeScript files in the dataset). Of the 100 most duplicated files, 98 are written in TypeScript. Collectively, these files account for 9.8 % of 3.9 million TypeScript files in the dataset. The most duplicated C files are mainly hardware drivers implementations.
0101020203030404050506060707080809090100100 Words (transcendental & misc. keywords) Words (all keywords) Files (transcendental & misc. keywords) Files (all keywords) Proportion of code (%)$C$ HeadersC++C++ HeadersJavaTypeScriptTypeScript HeadersC#Go
@@
We consider several metrics to identify which languages are most used for floating-point computations in our dataset. This comparison is illustrated in Fig.˜3. Each bar represents the proportion of files or words in the dataset that contain floating-point keywords, grouped by language. The first and third bars include all categories of keywords, while the others exclude keywords corresponding to floating-point types. We detail the choice of metrics and the corresponding results below.
  • Proportion of files with floating-point keywords (first bar): Counting the number of files, most of the code with any floating-point keyword is written in Java (20 %), while C++-specific headers are less common (2.2 %). The high number of Java files can be partly explained by the convention of isolating each class in its own file. In TypeScript, the use of the number type to represent both integers and floating-point values inflates the keyword count. To mitigate this effect, we also report the proportion of files excluding floating-point types and including only transcendental functions and miscellaneous keywords.
  • Proportion of files with transcendental functions and miscellaneous keywords (second bar): In this case, C++ source files make up most of the code (28 %). Notably, the proportion of TypeScript drops to 4.7 % for source files and 0.62 % for declaration files.
@@ 5.3. Per-Function Analysis
Based on the observation of the previous section, we parse source files written in C, C++, Java, TypeScript, C# and Go and extract functions containing floating-point keywords (steps 8 and 9 in Fig.˜1). For each function, we count the number of lines of code, words, parameters, floating-point keywords of the different categories, loops, conditional statements, function calls and maximum nesting depth of these constructs. This is the dataset we publicly release, which can be used to evaluate floating-point verification tools. Depending on the features a new verification tool is targeting, the developer can filter the dataset to sample relevant functions leveraging the above metrics.
[t] Lang # Filesa Files w/ fnb Files w/ fn w/ kwc # Fn w/ kwd C $515$ k 97 % 62 % 52 % $\pm$ 40 % C++ $888$ k 99 % 82 % 41 % $\pm$ 33 % C# $530$ k 89 % 73 % 41 % $\pm$ 32 % Go $136$ k 95 % 70 % 31 % $\pm$ 31 % Java $1.01$ M 93 % 82 % 43 % $\pm$ 34 % TS $706$ k 58 % 71 % 48 % $\pm$ 33 %
@@
Our final dataset consists of 10 105 03210\,105\,032 functions, of which 1 167 6531\,167\,653 are written in C (12 %), 4 272 6204\,272\,620 in C++ (42 %), 2 604 9192\,604\,919 in Java (26 %), 740 828740\,828 in TypeScript (7 %), 1 022 2661\,022\,266 in C# (10 %), and 296 746296\,746 in Go (3 %). Table˜4 reports on function complexity and the use of different language constructs in this dataset. For each construct, we report the proportion of functions that use it, along with the median and interquartile range (IQR, i.e. the difference between the 75th and 25th percentiles) of its occurrences among those functions. For each keyword category, we report the proportion of functions using at least one keyword from that category.
@@
In addition to the information in the table, we also analyze combinations of constructs: in C# and TypeScript, 11 % of functions include loops, conditionals, and function calls within a single function, compared to 26 % in C. When a function contains a loop, a conditional is also present in 69 % (Java) to 87 % (Go) of functions.
@@
Table 4. Function statistics per file. Med denotes median and IQR the interquartile range.
@@
Lang Words Params Loops Conditionals Function calls Keywords Med IQR Med IQR > 0 Med IQR > 0 Med IQR > 0 Med IQR Types Trasc Misc C 63 123 3 3 36 % 2 2 58 % 3 6 84 % 5 11 91 % 10 % 7.3 % C++ 43 95 2 2 22 % 2 2 44 % 3 5 76 % 6 14 95 % 6 % 5.7 % C# 34 63 2 2 15 % 1 1 42 % 2 4 73 % 3 7 97 % 5.7 % 3.5 % Go 42 97 2 2 29 % 2 2 63 % 2 4 85 % 6 14 97 % 3.4 % 5.2 % Java 27 61 1 2 20 % 1 1 37 % 2 3 73 % 5 10 98 % 4.6 % 3.9 % TS 27 57 2 1 15 % 1 1 44 % 2 3 83 % 3 6 98 % 1.5 % 4.0 % FPBench 23 26 2 1 12 % 2 2 6.9 % 1 0 61 % 1 1 100 % 38 % 0 %
@@
Answer to RQ2: In our dataset, conditional statements appear in many functions, up to 64 % in C. This proportion increases when loops are present, although loops themselves are less frequent. Function calls are also common, and typically these are not calls to transcendental functions. Special values like NaN or infinity appear rarely but are present. Despite these complexities, functions tend to be small, highlighting the importance of modularity in floating-point reasoning techniques.
@@
We compare our dataset to FPBench (Damouche et al., 2017), a benchmark suite of 131 floating-point functions written in Racket and designed for verification tools. FPBench functions can be exported to several supported languages. We export them to C and run our parser to extract statistics (see last row of Table˜4).
The FPBench functions have notably different characteristics: conditional statements and function calls appear less frequently, while calls to transcendental functions appear more often. The combination of a loop, conditional statement and a function call occurs only 1.5 % of the time (compared to 26 % in C). When a function contains a loop, a conditional is also present in 27 % of the cases (compared to over 69 % in real-world code). FPBench does align with our dataset in terms of the number of function parameters. The word count may be biased by the code being automatically exported to C from Racket and thus not written like a human would, e.g. with fewer intermediate variables.
@@
The results are summarized in Table˜5. We observe that GSL modules are not widely used in the dataset. Only 0.083 % of C files and 0.023 % of C++ files with floating-point keywords include functions from the GSL Math module. The Stats and SF modules appear even less frequently. In comparison, the math.h library is included in 16 % of C and 14 % of C++ files with floating-point keywords. Usage at the function level is similarly sparse as only 0.13 % of C floating-point functions include calls to GSL SF functions.
[t]   Module   Files   Functions   C   C++   C   C++   GSL Matha   0.083 %   0.023 %   0.072 %   0.0073 %   GSL SFb   0.070 %   0.028 %   0.13 %   0.011 %   GSL Statsc   0.022 %   0.0037 %   0.014 %   0.0017 %
@@
Answer to RQ3: FPBench functions are representative of real-world floating-point code in terms of the number of parameters but not in terms of complexity. Conditionals are rare, and function calls, when present, typically involve transcendental functions. Functions from the GNU Scientific Library that have been used for evaluation appear rarely in our dataset.

### v2

@@ 5.2. File-Based Analysis
After downloading the repositories and retaining only the files written in statically typed languages with floating-point keywords, we are left with 13 167 64813\,167\,648 files. To estimate the false positive rate, we draw a random sample from this dataset and manually inspect each file to determine whether it contains actual floating-point code. Using Cochran’s formula (cochran1977sampling) for sample size estimation in large populations, we select 385 files, ensuring statistical representativeness at a 95 % confidence level with a 5 % margin of error. Of these 385 files, 314 contain valid floating-point code, while 71 are false positives. Among the false positives, 50 arise from keywords appearing in comments, 9 from keywords in string literals, 7 from keywords used in variable names, and 5 from other miscellaneous causes. In most cases, the keyword double appears in comments in C and C++ files. This indicates the importance of stripping comments and ignoring string literals before performing keyword-based detection at the function level in later steps of the methodology.
Given the results of the manual inspection and using Wilson score intervals (Wilson01061927), we estimate that the proportion of true positives in our dataset lies between 77 % and 86 %, with 95 % confidence. Assuming that the presence of false positives in files is independent across files, the probability that project $i$ contains at least one file with actual floating-point code is
@@
That said, some files may have been missed due to the use of type-aliases or user-defined wrappers, so the numbers may be even higher. To assess the extent of this issue, we randomly sample 385 projects and manually inspect whether they redefine floating-point types. Among these, 22 projects do redefine floating-point types. Example of such aliases include declarations like type money = number in TypeScript, or conditional typedefs that map float and double to a common type Float to select precision at compile time with macros in C++. In all observed cases, these aliases are used alongside standard floating-point types rather than replacing them entirely, making the precise impact of this phenomenon difficult to quantify. An additional 10 projects define wrappers around floating-point values, typically simple structural wrappers in abstract syntax tree nodes. Over the full dataset, we therefore estimate with 95 % confidence that between 5.9 % and 11.5 % of projects are affected by floating-point type redefinitions or wrappers.
De-Duplication
A substantial fraction of the dataset is redundant, with 63 % of files identified as token-level duplicates and 61 % as exact duplicates. TypeScript exhibits the highest duplication rate, with declaration files duplicated on average 17 times and source files twice. C follows, with 3.4 copies per source file and 2.8 per header file. The most duplicated file originates from Angular, a TypeScript web application framework, and appears 27 13627\,136 times (0.70 % of TypeScript files in the dataset). Of the 100 most duplicated files, 98 are written in TypeScript. Collectively, these files account for 9.8 % of 3.9 million TypeScript files in the dataset. The most duplicated C files are mainly hardware drivers implementations.
0101020203030404050506060707080809090100100 Words (excluding type keywords) Words (all keywords) Files (excluding type keywords) Files (all keywords) Proportion of code (%) inCC HeadersC++C++ HeadersJavaTypeScriptTypeScript HeadersC#Go
@@
Stacked bars showing proportions of C, C++, Java, TypeScript, C#, and Go in files and words containing floating-point keywords. All bars go beyond 92% cumulatively, with C and C++ dominating.
@@
We consider several metrics to identify which languages are most used for floating-point computations in our dataset. This comparison is illustrated in Figure˜3. Each bar represents the proportion of files or words in the dataset that contain floating-point keywords, grouped by language. The first and third bars include all categories of keywords, while the others exclude keywords corresponding to floating-point types. We detail the choice of metrics and the corresponding results below.
  • Proportion of files with floating-point keywords (first bar): Counting the number of files, most of the code with any floating-point keyword is written in Java (20 %), while C++-specific headers are less common (2.2 %). The high number of Java files can be partly explained by the convention of isolating each class in its own file. In TypeScript, the use of the number type to represent both integers and floating-point values inflates the keyword count. To mitigate this effect, we also report the proportion of files excluding floating-point types and including only transcendental functions, arbitrary-precision and miscellaneous keywords.
  • Proportion of files with transcendental functions, arbitrary precision and miscellaneous keywords (second bar): In this case, C++ source files make up most of the code (28 %). Notably, the proportion of TypeScript drops to 4.6 % for source files and 0.62 % for declaration files.
@@ 5.3. Per-Function Analysis
Based on the observation of the previous section, we parse source files written in C, C++, Java, TypeScript, C# and Go and extract functions containing floating-point keywords (steps 8 and 9 in Figure˜1). We again estimate the false-positive rate by manually inspecting a representative sample of 385 extracted functions. Among these, 368 functions (96 %) contain valid floating-point code. The remaining false positives fall into two categories: nine cases in which floating-point keywords are used as variable or library names, and eight functions that include floating-point types only in parsing-related code. Based on this sample, we estimate with 95 % confidence that the true positive rate of our extraction lies between 93 % and 98 %.
For each function, we count the number of lines of code, words, parameters, floating-point keywords of the different categories, loops, conditional statements, function calls and maximum nesting depth of these constructs. This is the main dataset we publicly release. Depending on the features a new verification tool is targeting, the developer can filter the dataset to sample relevant functions leveraging the above metrics.
@@ C $508$ k 97 % 62 % 53 % $\pm$ 40 %
Our final dataset consists of 9 945 2539\,945\,253 functions, of which 1 168 2171\,168\,217 are written in C (12 %), 4 176 6354\,176\,635 in C++ (42 %), 2 574 3412\,574\,341 in Java (26 %), 723 609723\,609 in TypeScript (7 %), 1 009 3291\,009\,329 in C# (10 %), and 293 122293\,122 in Go (3.0 %). Table˜4(a) reports on function complexity and the use of different language constructs in this dataset. For each construct, we report the proportion of functions that use it, along with the median and interquartile range (IQR, i.e., the difference between the 75th and 25th percentiles) of its occurrences among those functions. For each keyword category, we report the proportion of functions using at least one keyword from that category. For illustration purposes, Figure˜4 shows a randomly selected function with a miscellaneous keyword (isFinite); note that no other floating-point related keywords appear.
@@
Table 4. Function statistics per function. FPB = FPBench. Med denotes median and IQR the interquartile range, computed out of positive occurrences only.
@@
Lang Words Params Loops Conditionals Function calls Keywords med IQR med IQR > 0 med IQR > 0 med IQR > 0 med IQR Types Trasc Misc MPFR C 61 121 3 4 36 % 2 2 57 % 3 6 84 % 5 12 91 % 10 % 7.1 % 0.38 % C++ 42 94 2 2 22 % 2 2 44 % 3 5 76 % 6 14 95 % 6 % 5.8 % 0.06 % C# 34 63 2 2 15 % 1 1 42 % 2 4 73 % 3 7 97 % 5.7 % 3.5 % 0.01 % Go 42 97 2 2 29 % 2 2 63 % 2 4 85 % 6 14 97 % 3.4 % 5.2 % 0.48 % Java 28 61 1 2 20 % 1 1 37 % 2 3 73 % 5 10 98 % 4.6 % 3.9 % 0.002 % TS 27 43 1 1 15 % 1 1 44 % 2 3 83 % 3 6 98 % 1.6 % 4.0 % FPB 23 26 2 2 12 % 1 0 6.9 % 1 1 61 % 2 1 100 % 38 % 0 % 0 %
(a) Floating-point code
Lang Words Params Loops Conditionals Function calls
med IQR med IQR > 0 med IQR > 0 med IQR > 0 med IQR
C 42 60 2 2 18 % 1 1 66 % 2 3 91 % 4 6
C++ 26 40 2 2 13 % 1 1 37 % 2 3 81 % 3 7
C# 23 27 1 2 5.5 % 1 1 29 % 2 2 74 % 2 3
Go 28 45 2 2 15 % 1 1 54 % 2 2 80 % 4 8
Java 13 25 1 1 9.9 % 1 1 24 % 1 2 61 % 3 6
TS 17 28 1 2 6.8 % 1 0 33 % 1 2 76 % 2 4
(b) All code
func writeSVG(writer io.Writer, f biFunc) {
fmt.Fprintf(writer, "<svg xmlns=’http://www.w3.org/2000/svg’ " +
"style=’stroke: grey; fill: white; stroke-width=0.7’ " +
$d’>\n", width, height$)
for i := 0; i < cells; i++ {
for j := 0; j< cells; j++ {
ax, ay := corner(i + 1, j, f); bx, by := corner(i, j, f)
cx, cy := corner(i, j + 1, f); dx, dy := corner(i + 1, j + 1, f)
if !isFinite(ax) $|$ !isFinite(ay) $|$ !isFinite(bx) $|$ !isFinite(by) $|$
!isFinite(cx) $|$ !isFinite(cy) $|$ !isFinite(dx) $|$ !isFinite(dy) {
continue
r, g, b := color(i, j, f)
fmt.Fprintf(writer, "<polygon points = ’%g,%g,%g,%g,%g,%g,%g,%g’
[…]
Figure 4. Randomly picked Go floating-point function in our dataset that contains miscellaneous keywords (writeSVG).
Go function example showing use of isFinite to filter non-finite floating-point values during SVG generation.
void RolloutStorage::compute_returns(torch::Tensor next, bool use_gae, float gamma, float tau) {
if (use_gae) {
value_predictions[-1] = next;
torch::Tensor gae = torch::zeros({rewards.size(1), 1}, torch::TensorOptions(device));
for (int step = rewards.size(0) - 1; step >= 0; --step) {
auto delta = rewards[step] + gamma * value_predictions[step + 1] * masks[step + 1] - value_predictions[step];
gae = delta + gamma * tau * masks[step + 1] * gae;
returns[step] = gae + value_predictions[step];
else {
returns[-1] = next;
for (int step = rewards.size(0) - 1; step >= 0; --step)
returns[step] = returns[step + 1] * gamma * masks[step + 1] + rewards[step];
Figure 5. Randomly picked C++ floating-point function in our dataset that contains both a loop and a conditional statement (compute_returns).
C++ function computing intermediate floating-point points along a line segment and returning them as cell coordinates.
In addition to the information in the table, we also analyse combinations of constructs: in C#, 10 % of functions include loops, conditionals, and function calls within a single function, compared to 26 % in C. When a function contains a loop, a conditional is also present in 69 % (Java) to 87 % (Go) of functions. Figure˜5 is an example of a randomly selected C++ function from our dataset that contains both a loop and a conditional statement.
We compare these results with those obtained from non-numerical code. We uniformly sample 10 00010\,000 projects from our dataset and apply the same methodology, yielding 9 708 1839\,708\,183 functions. Characteristics of these functions are summarised in Table˜4(b).
Compared to functions sampled from all projects, floating-point functions tend to be larger and contain loops more frequently. The larger average size of numerical functions may partly result from our keyword-based filtering, as larger functions are more likely to contain such keywords. The increased presence of loops could be a consequence of the larger average size of numerical functions. To assess whether function size alone accounts for the observed disparity in loop presence, we perform a controlled analysis. We fit a binary logistic regression model (hosmerAppliedLogisticRegression2013) predicting the presence of a loop as a function of the logarithm of function size and whether the function contains floating-points. After controlling for size, numerical functions remain significantly more likely to contain loops, with coefficients ranging from 0.1215 $\pm 0.003$ for C++ to 0.6395 $\pm 0.007$ for C#, all with $p<0.001$. According to the model, for functions of equal size, the odds of containing a loop are on average 13 % (C++) to 90 % (C#) higher in floating-point functions than in the general population. Although these results should be interpreted cautiously, they indicate that floating-point functions exhibit a systematically higher prevalence of loops beyond what can be explained by function size alone. The models achieve ROC AUC values between 0.79 (C) and 0.86 (Java). We do not observe a comparable effect for the presence of conditionals.
Mixed-Precision Code
We also analyse the use of different floating-point precisions within functions in our dataset. Table˜5 reports, for each precision, the proportion of functions in which a corresponding type keyword appears, as well as the proportion of functions that use keywords from multiple precisions within the same function. We omit TypeScript from this analysis, as it provides only a single floating-point type (number). For C and C++, we additionally report half precision (16-bit) and quad precision (128-bit), which are supported by these languages.
Single (32-bit) and double (64-bit) precision are the most commonly used across all languages, athough their relative prevalence varies substantially by language. This variation may be partly explained by differences in the libraries commonly used in each ecosystem. For example, in C#, where 65 % of functions use single precision, the Unity library predominantly relies on 32-bit floating-point data types such as Vector3 and Quaternion. In contrast, Go uses double-precision keywords in 82 % of functions, consistent with common practice in Go ecosystems where numerical libraries often operate on 64-bit floating-point values. Half (16-bit) and quad (128-bit) precision, as well as mixed-precision code, occur only rarely.
Table 5. Distribution of floating-point precision keywords used in functions per language: half (16-bit), single (32-bit), double (64-bit), quad (128-bit).
[t]
Module Precision Mixed-Precision
Half Single Double Quad
C 2.7 % 41 % 52 % 1.9 % 6.5 %
C ++ 0.1 % 60 % 38 % 0.7 % 3.9 %
C# 65 % 35 % 2.7 %
Go 23 % 82 % 8.5 %
Java 37 % 65 % 3.8 %
Parsing-Errors
Because not every source file is successfully parsed by Tree-sitter, we investigate how many files are affected by parsing errors and whether this biases our dataset toward parser-supported constructs. Tree-sitter recover from parsing errors by skipping invalid tokens and inserting placeholder nodes where syntactic elements are expected. Across the full dataset, Tree-sitter generate such error-recovery nodes in 16 % of the files. The languages most affected are C and C++, with 53 % and 33 % of files affected, respectively. Java and Go are significantly less affected, with 0.50 % and 0.23 % of files, respectively.
To assess the impact of these errors and characterise their causes, we manually inspect a sample of 97 files affected by parsing errors (corresponding to a Cochran sample size with a 10 % margin of error). In 90 out of 97 cases, parsing errors do not affect function extraction, i.e., no functions are skipped. Based on this observation, we estimate with 95 % confidence that between 0.57 % and 2.3 % of files in the full dataset have functions missed due to parsing errors. In most cases (74 out of 97), parsing errors stem from the use of preprocessor directives in C and C++ code. A common scenario involves conditional compilation directives (e.g., #ifdef), which may introduce unmatched or duplicated closing braces at the end of a function. Because preprocessor directives can only be resolved with full knowledge of the surrounding code, they are difficult to be handled reliably when analyzing files in isolation. Other sources of parsing errors include macros used as keywords (e.g., replacing static or inline) or macros that expand to language constructs such as for loops. We also observe cases in which Tree-sitter cannot correctly parse macro arguments without resolving macro expansion, for example when those arguments contain statements or type declarations. Finally, some errors (11 out of 97) arise from syntactic forms not supported by the Tree-sitter grammar, such as new ptr*[n].
@@
Answer to RQ2: In our dataset, conditional statements appear in many functions, up to 63 % in Go. This proportion increases when loops are present, although loops themselves are less frequent. For functions of equal size, floating-point functions are more likely to contain loops than functions sampled from all code. Function calls are also common in floating-point code, and typically these are not calls to transcendental functions. Special values like NaN or infinity appear rarely but are present. Despite these complexities, functions tend to be small, highlighting the importance of modularity in floating-point reasoning techniques. Single- and double-precision floating-point types are the most commonly used across all languages, whereas other precisions, arbitrary precision, and mixed-precision code are relatively rare.
@@
We compare our dataset to FPBench (Damouche et al., 2017), a benchmark suite of 131 floating-point functions written in Racket and designed for verification tools. FPBench functions can be exported to several supported languages. We export them to C and run our parser to extract statistics (see last row of Table˜4(a)).
The FPBench functions have notably different characteristics: conditional statements and function calls appear less frequently, while calls to transcendental functions appear more often. The combination of a loop, conditional statement and a function call occurs only 1.5 % of the time (compared to 26 % in C). When a function contains a loop, a conditional is also present in 27 % of the cases (compared to over 69 % in real-world code). FPBench does align with our dataset in terms of the number of function parameters. The word count may be biased by the code being automatically exported to C from Racket and thus not written like a human would, e.g., with fewer intermediate variables.
We presume that this difference arises because the FPBench benchmarks are drawn from existing literature and consist of hand-picked programs that current tools can already analyse successfully. In particular, many static floating-point reasoning tools offer limited support for conditionals, loops, and arbitrary function calls, which is reflected in the low prevalence of such constructs in FPBench.
@@
The results are summarised in Table˜6. We observe that GSL modules are not widely used in the dataset. Only 0.084 % of C files and 0.023 % of C++ files with floating-point keywords include functions from the GSL Math module. The Stats and SF modules appear even less frequently. In comparison, the math.h library is included in 17 % of C and 14 % of C++ files with floating-point keywords. Usage at the function level is similarly sparse as only 0.13 % of C floating-point functions include calls to GSL SF functions.
@@
Answer to RQ3: FPBench functions are representative of real-world floating-point code in terms of the number of parameters but not in terms of complexity. Conditionals are rare, and function calls, when present, typically involve transcendental functions that are over-represented. Functions from the GNU Scientific Library that have been used for evaluation appear rarely in our dataset.